OpenStudy (anonymous):

Rick has $1100 in the bank and is feeling pretty rich! He takes out $60 the first week ad continues to take out $60 every week. What equation will model the amount of money he has in the bank in any given week? Week 0 1 2 3 4 5 6 7 8 $ in the bank 1100 1040 980 920 860 800 740 680 620

OpenStudy (anonymous):

a=1100,d=60\[a _{n}=1100+(n-1)(-60)\]

OpenStudy (anonymous):

the general term can be written as \[a _{n-1}=1100+(n-1)(-60)\]